OpenStudy (anonymous):

Consider the following 3x4 contingency table in which a sample of 1400 companies is summarized in terms of the company's industry type (X, Y, or Z) and geographic location (A, B, C, or D). A B C D X 124 110 108 114 Y 130 106 122 112 Z 128 118 112 116 What is the probability P(B∪Y)?

OpenStudy (anonymous):

P(B) = (110+106+118)/1400 P(Y) = (130+106+122+112)/1400 P(B∪Y) = P(B) + P(Y) - 106

OpenStudy (anonymous):

erm, make that P(B) + P(Y) - 106/1400